OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for x in the first equation. x + 5y - 10 = 0 x = 2y - 8 What is the resulting equation in simplest form? -3y - 18 = 0 3y - 18 = 0 7y - 18 = 0 i got a

OpenStudy (opcode):

Mr. Dan give me a moment I work better with paper and pencil on these types of questions :-).

OpenStudy (anonymous):

thanks and ok

OpenStudy (opcode):

I take this is Algebra II? Anyway here is my work: x + 5y - 10 = 0 x = 2y - 8 2y - 8 + 5y - 10 = 0 7y - 18 = 0 y = 2.57142857143 (I'll prettify this with LaTeX in a second.)

OpenStudy (anonymous):

C

OpenStudy (opcode):

I take this is Algebra II? Anyway here is my work: First equation: \[x+5y−10=0\] Second equation: \[x=2y−8\] They want us to use the above equation to get a solution for: \(x+5y−10=0 \) Input \(2y−8\) for \(x\) So: \[2y−8+5y−10=0\] \[2y+5y=7y\] \[−8−10=−18\] So the equation in the simplest form is: \[7y−18=0\]